### 25 Jan 2024 v1

# Knowledge Graph Supported Benchmark and Video Captioning for Basketball

Abstract
Despite the recent emergence of video captioning models, how to generate the text description with specific entity names and fine-grained actions is far from being solved, which however has great applications such as basketball live text broadcast. In this paper, a new multimodal knowledge supported basketball benchmark for video captioning is proposed. Specifically, we construct a Multimodal Basketball Game Knowledge Graph (MbgKG) to provide knowledge beyond videos. Then, a Multimodal Basketball Game Video Captioning (MbgVC) dataset that contains 9 types of fine-grained shooting events and 286 players’ knowledge (i.e., images and names) is constructed based on MbgKG. We develop a novel framework in the encoder-decoder form named Entity-Aware Captioner (EAC) for basketball live text broadcast. The temporal information in video is encoded by introducing the bi-directional GRU (Bi-GRU) module. And the multi-head self-attention module is utilized to model the relationships among the players and select the key players. Besides, we propose a new performance evaluation metric named Game Description Score (GDS), which measures not only the linguistic performance but also the accuracy of the names prediction. Extensive experiments on MbgVC dataset demonstrate that EAC effectively leverages external knowledge and outperforms advanced video captioning models. The proposed benchmark and corresponding codes will be publicly available soon.
1 Introduction
Video captioning (VC) is a crucial computer vision task that requires the model to outputs corresponding text descriptions based on the given video. Despite the recent rapid development of video captioning [17, 12, 16, 13, 36, 33], how to automatically generate a text description with specific entity names and fine-grained actions is seldom investigated. As the basketball live text broadcast example in Fig. 1, the basketball video involves multi-person actions and complex scenes, which pose significant challenges to model’s performance and generalization. Note that conventional models can only generate the simple sentence to describe the video from a macroscopic perspective (e.g., a man fails to make a shot and another man gets the rebound). In contrast, if the model has game-related knowledge such as players who appear in the game and fine-grained actions in basketball, it can generate the knowledge-grounded text description (e.g., Brandon Ingram misses the 2pt jump shot and Justise Winslow gets the defensive rebound).
Figure 1: Comparison of conventional captioning with knowledge-grounded captioning. The different specific entity names are marked red and blue, respectively. And the fine-grained actions are marked green.
In addition, existing common used benchmarks, including MSVD [2], YouCook [3], MSR-VTT [31], and ActivityNet Captions [9], simplify the task by using indefinite pronouns like “a man”, “a woman” or “a group of men” instead of specific entity names. And the actions in the annotated captions are coarse-grained. These benchmarks cannot provide relevant knowledge beyond videos and fine-grained action annotations to develop models to generate text descriptions with specific entity names and fine-grained actions. To our best knowledge, there is no relevant benchmark in basketball that can support this task.
Figure 2: An example of a Multimodal Basketball Game Knowledge Graph.
To narrow this gap, a new multimodal knowledge graph supported basketball benchmark for video captioning is proposed. This benchmark requires the model to comprehend the video’s content and generate event descriptions that incorporate specific entities with additional knowledge. Specifically, we collect 25 NBA full-games multimodal data from professional basketball platforms in 2022-2023 season, including events, players’ information, teams’ information, and videos. Based on the collected data, we construct a Multimodal Basketball Game Knowledge Graph (MbgKG) to provide the knowledge beyond videos, as depicted in Fig. 2. Subsequently, a Multimodal Basketball Game Video Captioning (MbgVC) dataset is constructed based on nodes in MbgKG and relationships among the selected nodes. MbgVC dataset comprises 9 types of basketball shooting events and 286 players’ knowledge (i.e., images and names), with data samples illustrated in Fig. 3.
To effectively utilize knowledge in the MbgVC dataset, we propose an encoder-decoder framework named Entity-Aware Captioner (EAC) for basketball live text broadcast. EAC is comprised of 3 separate unimodal encoders for videos, players’ images, and players’ names, as well as a pre-trained language model that serves as the text decoder for generating descriptions of the given videos. Moreover, the Bi-GRU module is introduced to encode the temporal information, while the multi-head self-attention module is introduced to model the associations among the candidate players and select the key players. In addition, we propose a new performance evaluation metric, the Game Description Score (GDS), which takes into account both the linguistic quality of the description (as used in conventional video captioning tasks) and the accuracy of name predictions.
The main contributions of this paper are as follows:
  • A Multimodal Basketball Game Knowledge Graph (MbgKG) is constructed to provide additional contextual knowledge that extends beyond the video contents.
  • Based on MbgKG, nodes and relationships are extracted to construct our Multimodal Basketball Game Video Captioning (MbgVC) benchmark. Additionally, we propose a new evaluation metric for MbgVC that offers a more accurate and comprehensive evaluation by incorporating the accuracy of name predictions.
  • We develop an Entity-aware Video Captioning framework in encoder-decoder form, which selectly incorporates key information from external knowledge to generate text descriptions with specific entity names.
  • On the MbgVC, our proposed model outperforms existing advanced models, achieving leading performance.
Figure 3: Data samples from the proposed dataset. Each video is annotated by fileid, action type, caption, player images and player names. Each of the players involved in caption as well as their teammates serve as candidate players.
2 Related Works
2.1 Computer Vision with Knowledge Graph
A knowledge graph (KG) is essentially a large-scale semantic network that contains entities, concepts as nodes and various semantic relationships among them as edges [37]. The application of knowledge graphs have greatly promoted the rapid development of computer vision. Zhuo et al. [38] utilize the mined category-attribute relationships in knowledge graph and the similarity between the seen categories and the unseen ones for more reliable knowledge transfer. One of the main challenges in video-text retrieval task is to identify fine-grained semantic associations between video and text. To address this issue, under the guidance of external knowledge, Fang et al. [4] utilize the associations between concepts to extend new concepts and enrich the representation of videos. This approach enables a more accurate matching of video and text, leading to improved retrieval results. Gu et al. [6] propose a text with knowledge graph augmented transformer, which integrates external knowledge and leverages multimodal information to mitigate the challenges posed by long-tail words. However, unlike the aforementioned works that extract knowledge from unimodal knowledge graphs, our approach extracts knowledge from a multimodal knowledge graph, utilizing this multimodal knowledge to assist the model in generating descriptions that include specific entity names and fine-grained actions.
2.2 Video Captioning Benchmarks
Existing widely used benchmarks to support the video captioning task include MSVD [2], MSR-VTT [31], YouTube Hightlight [25] and VATEX [30]. These benchmarks are open-domain databases that enhance the generalization of models. However, these benchmarks suffer from common limitations: the text descriptions are too concise, ignoring the names of specific entities and fine-grained action types. Some works also attempt to enrich benchmarks and develop models to generate text descriptions with more fine-grained information. Fang et al. [5] construct a benchmark annotated with captions and commonsense descriptions. This commonsense benchmark develops models to generate captions as well as 3 types of commonsense descriptions (intention, effect, and attribute). Yu et al. [34] propose a fine-grained sports narrative benchmark that focuses more on the detailed actions of the subjects, but this benchmark ignores the names of specific subjects. To assist visually impaired individuals in enjoying movies, Yue et al. [35] construct a large-scale Chinese movie benchmark, which requires models to generate role-aware narration paragraphs when there are no actors speaking. Qi et al. [19] utilized a unimodal knowledge graph for real-time soccer commentary, incorporating all pre-match and player-team information without a selection mechanism. This indiscriminate inclusion could introduce noise, affecting caption accuracy. Contrastingly, our approach first employs a multimodal knowledge graph, integrating both textual and visual information, such as player names and images. This enriches the context for video captioning, improving the match with video content. Second, we propose a selective attention mechanism, modeling associations among candidate players and incorporating key players’ knowledge. This effectively reduces noise and enhances caption precision. These advancements distinguish our work from Qi et al.’s [19] and other existing models, demonstrating the potential of multimodal knowledge and selective attention in video captioning tasks.
2.3 Video Captioning Models
Figure 4: Illustration of the proposed video captioning model that leverages external knowledge from multimodal knowledge graph.
Video captioning is a crucial task in video understanding, where models generate text descriptions for given videos. Some works [10, 14, 16, 36] utilize a visual encoder to extract video representations from a set of video frames, and a language decoder transfer these representations to the corresponding descriptions. Recently, CLIP (Contrastive Language-Image Pretraining) [20] has demonstrated its superior performance on various visual-linguistic tasks relying on large-scale contrastive pre-training with image-text pairs. Therefore, Clip4Caption [26] employs CLIP to acquire the aligned visual-text representation, leading to significant improvements in video captioning performance. Benefiting from the flexibility of the transformer [27] architecture, SwinBert [12] introduces Video Swin Transformer [15] as the video encoder to encode spatial-temporal representations from video frames. However, none of the previously mentioned works can address the challenge of generating action entity names effectively. In stark contrast, we propose to incorporate external knowledge based on an encoder-decoder model structure to generate text descriptions with specific entity names and fine-grained actions for basketball live text broadcasts, demonstrating the practicality of our entity-aware video captioning in real applications.
3 Proposed Benchmark
In this section, we first construct a multimodal knowledge graph as the knowledge base, and then extract relevant data and knowledge to construct a multimodal basketball dataset. Furthermore, a new metric for this dataset is proposed to assess the performance of models.
3.1 Multimodal Basketball Game Knowledge Graph
Our knowledge graph is based on basketball game videos with corresponding event descriptions. To begin the construction of the Multimodal Basketball Game Knowledge Graph, we collect full-game play-by-play data from 50 games, including descriptions of events, the time on the scoreboard corresponding to the events, score records, players’ information, and teams’ information, from the professional basketball data platform [^1] https://www.basketball-reference.com. The corresponding videos are collected from the basketball broadcast platform [^2] https://fishkernba.com. After filtering out videos with low resolution and chaotic content, only 25 games are retained.
To structure the collected data, the following steps are needed to process it: (1) Categorize basketball events. (2) Parse and structure the descriptions of each type of event. (3) Match these event descriptions with the corresponding video timestamps using OCR toolkits.
Labels
2p
-succ.
2p
-fail.-off
2p
-fail.-def.
2p
-layup
-succ.
2p
-layup
-fail.-off.
2p
-layup
-fail.-def.
3p
-succ.
3p
-fail.-off.
3p
-fail.-def.
Train num. 469 146 397 442 133 251 470 202 652
Test num. 95 41 95 108 32 67 125 61 162
Table 1: Statistics of the labels in NBA. “succ.” and “fail.” are abbreviations of “success” and “failure”, respectively.
By analyzing the play-by-play data, game events can be divided into 9 categories, including “Foul”, “Rebound”, “Violation”, “Timeout”, “Freethrow”, “Enter game”, “Turnover”, “Jump ball”, and “Shot”. Each text description has its owns keyword, such as the description of the “Foul” event “Personal foul by G. Temple (drawn by A. Drummond)”, where the keyword is “foul”. Therefore, the type of this event is “Foul”. For this semi-structured data, the sentence is parsed into multi-tuples using the character index of the keyword and specific words that appear in the sentence, such as “drawn by”. To incorporate video information into the multi-tuples, the video timestamp is intended to be associated with the text description. OCR is used to identify the time on the scoreboard in each frame and record the timestamp of the current frame. Specifically, we employ the open-source OCR toolkit PaddleOCR [^3] https://github.com/PaddlePaddle/PaddleOCR and Tesseract-OCR [24] to recognize the time simultaneously. By using multiple OCR toolkits, we can improve the overall accuracy of recognition. This is because one toolkit may succeed where the other fails. In addition to event information, player and team information are also stored in multi-tuples. Through these efforts, a Multimodal Basketball Game Knowledge Graph containing 11489 events and 42870 relationships is constructed. An example of the Multimodal Basketball Game Knowledge Graph (MbgKG) is shown in Fig. 2. More details on the construction of MbgKG can be found in Appendix A.
3.2 Multimodal Basketball Video Game Captioning Dataset
To construct a dataset that contains videos, text descriptions, and external knowledge (i.e., player images and names), the pertinent data is extracted from MbgKG by utilizing the relations within MbgKG. As shot and rebound events are the most prevalent in basketball, we only extract the data about these events from MbgKG. To be more realistic, the text descriptions need to be further modified.
When a shot fails to score in basketball, it is often followed by a rebound event. In line with this pattern, we have combined the shot and rebound into a single event. Shots in basketball can be categorized as two-point (2pt) shots, three-point (3pt) shots, and layups. Rebounds are further divided into defensive (def.) rebounds and offensive (off.) rebounds. The classification scheme of our dataset is based on the NBA dataset [32] used for group activity recognition. However, unlike the NBA dataset, where all videos are 6 seconds long and divided into 72 frames, we consider that different events may have different durations. Consequently, in our dataset, videos of different lengths are uniformly divided into 72 frames. Additionally, we have incorporated text descriptions and player information into our dataset.
As shown in Fig. 4, the event-action relation in the graph is used to extract 9 kinds of events, and the video-time relation in the graph is used to obtain the video timestamp of the event. Specifically, we first roughly estimate the start and end timestamps of the event based on the existing timestamp. Each clip undergoes a manual review process to ensure accurate start and end timestamps. Subsequently, the video-description relation is used to obtain the text description, which is then matched with the corresponding video clip. We extract player-related information (player images and player names) from MbgKG through the team-player and player-name relations. Since each event involves certain players, only the involved individuals and their teammates need to be considered as candidate players for video annotation.
Using the sentence “B. Ingram misses 2-pt jump shot from 19 ft and defensive rebound by J. Winslow” as an example, we will explain our text modification process. Due to the difficulty in generating distance in video captioning task, the “from 19 ft” in the sentence is removed. Based on the attribute information of the player in MbgKG, the abbreviated name (“J. Winslow”) in the sentence is replaced with the full name (“Justise Winslow”). This change might increase the complexity of name generation but is more aligned with real-world scenarios. To improve text fluency, alterations are made based on the character index. Specifically, we identify the index of keywords like ”defensive” and apply a rule (“defensive/offensive rebound by SOMEONE” $\Longrightarrow$ “SOMEONE gets the defensive/offensive rebound”) to revise the sentence. Finally, there are a total 3977 videos [^4] The scale of the current dataset is not large enough, which is emergent to be enriched in our subsequent research., each of which belongs to one of 9 kinds of events. Each video has one text description and several candidate players information (images and names). We randomly select 3162 clips for training and 786 clips for testing. Tab. 1 shows the sample distributions across different labels of events. More details on the construction of MbgVC can be found in Appendix B.
In Tab. 2, the comparison of our dataset with MSR-VTT, YouCook and ActivityNet Captions further demonstrates the fine-grained details of our captioning annotations. MbgVC has the most sentences per second of 0.182, while the other datasets are all below 0.1, this reflects that our dataset has more detailed information in descriptions. Moreover, MbgVC has 1.73 verbs in a sentence on average, higher than 1.41 for ActivityNet Captioning and 1.37 for MSR-VTT. Similarly, the verb ratio of MbgVC which is computed by dividing the total number of verbs by the total number of words in the sentence is also much higher than other 3 datasets. This highlights that our dataset is primarily focused on the fine-grained actions of the subjects, aligning with our original intent.
3.3 Evaluation Metric
Existing video captioning benchmarks mainly adopt n-gram based metrics, including CIDEr [28], METEOR [1], Rouge-L [11], and BLEU [18]. BLEU evaluates the quality of the generated text by calculating the n-gram overlap between the generated text and the reference text. ROUGE-L pays more attention to generating the longest common subsequence between the generated text and the reference text, and evaluates sentence quality in this way. METEOR takes more factors into account, including stem matching, synonym matching, and word order. CIDEr uses TF-IDF [22] to assign different weights to n-grams of different lengths, then calculates the cosine similarities of n-grams between generated texts and reference texts and averages them to get the final score. These metrics primarily emphasize the consistency of text rather than the accuracy of semantics. In other words, if the generated text differs from the reference in structure or vocabulary, it may be penalized by these n-gram based evaluation metrics, even if the conveyed information is completely correct. However, for entity names, they cannot be replaced with the similar words.
To take into account the predictive accuracy of specific entity names, the entity name prediction accuracy $e_{acc}$ is introduced to accurately evaluate the accuracy of name prediction. $e_{acc}$ is a score calculated by dividing the number of correctly predicted names by the total number of names. Finally, we settle on a new metric, Game Description Score (GDS), as follows:
$gds=\frac{4\cdot e_{acc}+2\cdot cs+2\cdot ms+1\cdot rs}{9}\times 100,$ (1)
where $cs$, $ms$, and $rs$ refer to entity name prediction accuracy score, CIDEr score, METEOR score, and Rouge-L score, respectively. Since CIDEr involves the partitioning and matching of n-grams, BLEU is not included in the calculation. More detailed descriptions of GDS can be found in Appendix C.
Dataset
Sentences
per second
Verbs per
sentence
Verb
ratio
MSR-VTT 0.067 1.37 14.8%
YouCook 0.056 1.33 12.5%
ActivityNet
Captions
0.028 1.41 10.4%
MbgVC 0.182 1.73 14.8%
Table 2: Comparisons of different video caption datasets. Numbers in bold indicate the highest value.
4 Knowledge Graph Supported Basketball Video Captioning
4.1 Problem Formulation
To generate precise and concise textual descriptions of the basketball live text broadcast, we introduce the concept of a Knowledge-Aided Basketball Video Captioning (KABVC) task. In this task, the model is required to comprehend the content of the provided video and generate text descriptions that include specific entity names and fine-grained actions based on the external knowledge. This task can be formulated as: given the basketball video $V_{b}$, the objective is to select video-related player knowledge $K_{p}=s\left(V_{b}\right)$ and generate the text description $D_{v,k}$.
${\rm KABVC}:D_{v,k}=m\left(V_{b},K_{p}\right)=m\left(V_{b},s\left(V_{b}\right)\right),$ (2)
where $s\left(\cdot\right)$ and $m\left(\cdot\right)$ denote the models’ abilities on aligning video and knowledge (player information) and transferring video and knowledge to the text description, respectively.
Model CIDEr METEOR Rouge-L BLEU-1 BLEU-2 BLEU-3 BLEU-4 GDS
V2C 13.7 18.7 45.9 50.1 39.5 27.0 14.9 12.3
Clip4Caption 19.4 26.7 51.2 49.1 42.5 35.4 28.8 25.0
SwinBert 20.1 26.5 49.0 47.8 41.4 34.5 28.4 25.6
EAC 138.5 28.0 54.9 53.1 46.4 38.8 32.4 56.9
Table 3: Combined inference performance on MbgVC. Numbers in bold indicate the best performance.
4.2 Proposed Model
For the ${\rm KABVC}$ task, we propose a model with an encoder-decoder framework named Entity-Aware Captioner (EAC). The overall structure of EAC framework is shown in Fig. 4. Given the raw video frames which are of size $T\times 3\times H\times W$, consisting of $T$ frames and each has $3\times H\times W$ pixels. We feed them into Resnet-18 [7], which is the video encoder in EAC, and extract the video feature $F_{T}\in{\rm R}^{T\times D_{r}\times h\times w}$. $D_{r}$ is the hidden size of Resnet-18. The global representations of $F_{T}$ are then fed into Bi-GRU module to further encoder temporal information and obtain the feature $F_{v}$.
$F_{v}=W_{1}\left({{\rm GRU}\left({\rm AvgPool}\left(F_{T}\right)\right)}\right),$ (3)
where $W_{1}\in{\rm R}^{D_{r}\times D}$ is the linear mapping layer, which maps the video feature to text space. $D$ is the hidden size of the decoder module. ${\rm GRU}\left(\cdot\right)$ denotes the Bi-GRU module and ${\rm AvgPool}\left(\cdot\right)$ denotes the average pooling layer.
Each video has the corresponding candidate $N$ players’ information as the external knowledge to assist the generation of text with the specific entity names. Each player image is of size $3\times H_{p}\times W_{p}$. For candidate players’ images, Resnet-18 is used to extract the features $F_{N}\in{\rm R}^{N\times D_{r}\times h_{p}\times w_{p}}$. The global features $F_{p}\in{\rm R}^{N\times D_{r}}$ of $N$ images are obtained by Eq. 4.
$F_{p}=W_{2}\left({\rm AvgPool}\left(F_{N}\right)\right)),$ (4)
where $W_{2}\in{\rm R}^{D_{r}\times D}$ is the linear mapping layer, which maps the image feature to text space.
The text encoder of large language model T5 [21] is used to transform candidate names into a sequence of embeddings $T_{n}\in{\rm R}^{N\times D}$. After that, the image features $F_{p}$ and name features $T_{n}$ are added by corresponding positions to obtain multimodal player features $F_{m}\in{\rm R}^{N\times D}$. Basketball events involve interactions between multiple players. For example, in “Myles Turner makes the 2pt layup with an assist from Tyrese Haliburton”, “Myles Turner” scores a two-point layup with the assist of “Tyrese Haliburton”. Therefore, the multimodal player features $F_{m}$ are fed into the multi-head self-attention module to model the relationship among candidate players. The player features based on attention and the video features containing motion information are then concatenated and fed into the decoder ${\rm\Psi_{T5}}\left(\cdot\right)$ of T5 to obtain the text description $C$.
$C={\rm\Psi_{T5}}\left(W_{3}\left({\rm Concat}\left[{\rm H_{M-ATT}}\left(F_{m}\right),F_{v}\right]\right)\right),$ (5)
where $W_{3}\in{\rm R}^{D\times D}$ is the linear mapping layer, which maps the concatenated feature to the vector space of the T5 model. ${\rm Concat}\left[,\right]$ denotes the Concatenation function in Python, and ${\rm H_{M-ATT}\left(\cdot\right)}$ denotes the multi-head self-attention module.
The Bi-GRU module, 3 linear layers, multi-head self-attention and text decoder are trained to maximize the log-likelihood over the training set given by Eq. 6.
$L_{\Theta}=\sum_{t=1}^{N_{C}}log{\rm P_{r}}\left(y_{t}|y_{t-1},\left[F_{m},F_{v}\right];\Theta\right),$ (6)
where $y_{t}$ denotes the one-hot vector probability of each word at time $t$. $N_{C}$ denotes the length of the caption. And $\Theta$ denotes the learnable parameters.
5 Experiments
To verify the performance of the proposed model, EAC is compared with the advanced video captioning models. We further conduct the ablation experiments to verify the effectiveness of each conponent in EAC.
5.1 Implementation Details
The proposed EAC uses ResNet-18 pretrained on the ImageNet dataset [23] as the visual encoder, which hidden size $D_{r}$ is 512. EAC is trained on the proposed MbgVC dataset with 100 training epochs. In MbgVC dataset, $T$ frames are sampled by using segment-based method [29]. Each frame size is $720\times 1280$. Concretely, $T$ is set to 18. The size of player’s image is $120\times 180$ and the number of candidate players in each video is not fixed. For the large language model T5, the hidden size is 768. It is worth noting that the parameters of the visual encoder and text encoder are frozen during training. During the training stage, the model is optimized by ADAM [8] with the learning rate of 3e-5 and weight decay of 1e-4. Beam search with a beam size of 2 is used for inference. The proposed EAC is implemented with Python 3.9 and PyTorch 1.12, and is performed on a server with an Nvidia 3090ti GPU.
5.2 Main Results
EAC is compared with three advanced video captioning models, including V2C [5], Clip4Caption [26], and SwinBert [12]. V2C is the Transformer-based model that generates relevant commonsense descriptions of the given video. Clip4Caption employs CLIP to acquire the aligned visual-text representation for better generating the text descriptions. SwinBert introduces Video Swin Transformer to encode spatial-temporal representations from video frames.
As shown in Tab. 3, EAC outperforms other 3 models by a large margin on CIDEr. This notable performance can be attributed to the way CIDEr calculates cosine similarities of n-grams between the generated text and reference text. Highly accurate name prediction contributes to a higher cosine similarity between the n-grams of the generated text and the reference text. V2C, Clip4Caption, and SwinBert exhibit lower performance in accurately generating names, resulting in much lower CIDEr scores. Although EAC achieves the best performance across all metrics, the differences are not as pronounced in METEOR, Rouge-L, and BLEU. This can be attributed to the concise nature of the text descriptions in the dataset, primarily consisting of names and fine-grained actions. Notably, EAC outperforms SwinBert 31.3% on GDS and outperforms Clip4Caption 31.9% on GDS. The higher GDS indicates better overall model performance, particularly in name prediction. These results underscore the capability of our model to generate accurate names and fine-grained actions in live text broadcast task. Qualitative results can be found in Appendix D.
Model Ki Kn Attn. Bi-GRU CIDEr METEOR Rouge-L BLEU-4 GDS
20.1 20.6 40.9 23.0 19.1
18.0 20.0 38.0 23.2 16.3
110.6 25.2 49.1 27.3 46.7
115.7 25.6 53.1 28.1 49.8
119.3 26.8 53.3 29.5 51.4
122.2 27.3 54.6 30.5 52.5
138.5 28.0 54.9 32.4 56.9
Table 4: Ablation study on MbgVC. Ki and Kn denote the player images knowledge and player names knowledge, respectively. Attn. is the multi-head self-attention module. Bi-GRU is bi-directional GRU module. Numbers in bold indicate the best performance.
5.3 Ablation Study
To verify the contribution of the external knowledge and other modules in EAC model, we also perform an ablation study by progressively adding these as input. The results of ablation study are shown in Tab. 4. Model ① consists of Resnet-18 and T5 decoder, with its input being solely video. From the comparison results between of model ① and model ②, it can be observed that adding players’ images to model based on ① does not improve the model’s performance. During the training stage, if the model has not been exposed to enough specific names, or has not learned how to infer names from video and image features, it may not be able to generate these names. The comparison results between model ① and model ③ highlight a significant improvement by solely adding players’ names to the model based on ①. Providing a list of names as input allows the model to effectively integrate this information with the video features, leading to more accurate generation of names. This is likely providing the model with external contextual information to help it make more accurate predictions. From the comparison results between model ③ and model ④, we can find that player images can emphasize the roles based on entity names and improve the names’ prediction. From the comparison results of models ④-⑦, it can be observed that adding multi-head self-attention and Bi-GRU modules can improve the performance of the model. The multi-head self-attention module can model the associations among players and focus on key players. The Bi-GRU module can model the temporal information of video frame features, capturing the action information. From the above results, the external players knowledge brings significant gains in entity awareness. The model can generate text descriptions with specific player names and fine-grained actions through external knowledge and temporal modeling.
5.4 Error Analysis
The error analysis are conducted by presenting cases of EAC. As shown in Fig. 5, we show several errors. (1) Player Mismatching: when the number of players in the sentence is more than one, the model may decode the name of one of the players incorrectly. (2) Action Confusion: model tends to confuse similar-looking actions, such as layups and close-range two-point jump shots. (3) Lack Distance-aware Perception: model tends to confuse three-point jump shots and long-range two-point jump shots. For example, the player appears to be shooting a three-point shot, but is actually inside the three-point line.
Figure 5: Representative error cases of the generated captions, which correspond to the player mismatching, action confusion and lack distance-aware perception. GT denotes the ground truth description and PR denotes the generated description.
Given the above result, we discuss some potential ways for developing an advanced models for KABVC task. First, beyond the current object detection, it is necessary to enhance the model’s ability to understand and handle names. Second, regarding action confusion, this may be because the model is not sensitive enough to subtle differences in the video to accurately distinguish similar actions. We need to improve the feature extraction abilities of the model to more accurately recognize similar-looking actions. Third, for the lack of distance perception, this may be because the model has trouble processing spatial information, especially when judging the relative distance of players from the basket. This indicates that we need to further improve the spatial awareness of our model.
6 Conclusion
In this paper, we introduce the task of knowledge aided basketball video captioning (KABVC) for live text broadcast. To achieve this task, we further construct a multimodal basketball game knowledge graph (MbgKG) to provide the external knowledge. Through partial relationships in the knowledge graph, a new multimodal basketball game video captioning dataset is then constructed. We also propose a new metric GDS for video captioning, which consider the accuracy of entity name prediction. Furthermore, our experiments validate the importance of additional knowledge, including player images and entity names, for basketball video captioning.
Appendix A Construction of MbgKG
Figure 6: An example of the play-by-play page.
Figure 7: Examples of 9 types of basketball events. Each type has the corresponding keyword marked in red.
Figure 8: Examples of player information and team information.
Figure 9: An example of the multi-tuples parsed from event descriptions, players’ information, and teams’ information. In this multi-tuples, the home team information is marked in red, the away team information is marked in blue, and events are marked in green.
The example of play-by-play page is shown in Fig. 6. This page not only contains the event descriptions, but also contains links to player information and team information. We can collect relevant players’ information and teams’ information through these links. All the collected event descriptions can be divided into 9 categories, as shown in Fig. 7. Each event has its own unique keywords which are marked in red. In addition to keywords, there are also some special words in different events. For example, in shooting events, there are words such as “assist by” and “block by”. By obtaining the indices of these keywords and special words, we can parse the descriptions into multi-tuples. Fig. 8 shows the players’ information and teams’ information. We only collect information that is available for most players, including “name”, “position”, “shoots”, “height”, “weight”, “birth”, “NBA debut”, and the link to their profile pictures. For teams’ information, we only collect “name”, “season”, “coach”, “executive”, “arena”, and the link to teams’ logos. By using OCR technology to recognize the time on the scoreboard in each video frame, we can compare it to the time on the play-by-play page. If the recognized time matches the time on the scoreboard, the video timestamp of the current event is recorded in the multi-tuples. Fig. 9 shows the example of the multi-tuples parsed from event descriptions, players’ information, and teams’ information. Through these efforts, a Multimodal Basketball Game Knowledge Graph (MbgKG) containing 11489 events and 42870 relationships is constructed. MbgKG not only contains the static information, but also includes the temporal dynamic information. We believe that this knowledge graph will continue to be extended and be explored more valuable knowledge for sports video understanding.
Appendix B Construction of MbgVC
In basketball, shot and rebound events are the most prevalent. Meanwhile, we count the number of different types of events in MbgKG. Tab. 5 shows that there are 4400 shot events and 2649 rebound events, which are the most common types of events. Based on the aforementioned common knowledge and statistical data, videos and event descriptions about shot and rebound are extracted from MbgKG to construct the dataset.
Shots can be divided into three types: two-point jump shot (2p jump shot), two-point layup (2pt-layup), and three-point jump shot (3p jump shot). Rebounds can be divided into two types: defensive rebound and offensive rebound. In addition, scoring and missing a shot are represented by “make” and “miss”, respectively. When a shot is not scored, it is often followed by a rebound event. Therefore, we combine the missed shot event and rebound event into one event. In total, there are nine types of shooting events, which are “2p-succ.”, “2p-fail.-off.”, “2p-fail.-def.”, “2p-layup-succ.”, “2p-layup-fail.-off.”, “2p-layup-fail.-def.”, “3p-succ.”, “3p-fail.-off.” and “3p-fail.-def.” Subsequently, based on these nine types of shooting events and the relationships within MbgKG, we extract videos, text descriptions, and player information from MbgKG. It’s worth noting that there are some issues with the text descriptions in MbgKG: 1) The names of the players are abbreviated, which does not meet the requirements of actual applications; 2) The texts contain descriptions of shooting distances, such as ”from xx ft” and ”at rim”. This makes it difficult for the captioning model to generate accurate distances; 3) The text descriptions do not conform to the structural and compositional rules of natural language.
Figure 10: Qualitative results on MbgVC dataset. (V2C: Video2Commonsense; C4C: Clip4Caption; SwB: SwinBert; EAC: our proposed model; GT: the ground truth). The different specific entity names are marked in red and blue, respectively. And the fine-grained actions are marked green. Since V2C does not have its own tokenizer and vocabulary list, it cannot decode names. So names are replaced with the special token $<$UNK$>$.
In response to the above issues, we design some modification rules to address them. For names, we replace abbreviated names in the text descriptions with the full names from the player attribute information in MbgKG (“L. James” $\Longrightarrow$ “LeBron James”). The distance descriptions (“from XX ft” and “at rim”) in the texts are directly removed. Sometimes, when a player scores a shot, it is assisted by his teammate. We modify “(assist by SOMEONE)” to “with an assist from SOMEONE”. When a player doesn’t score a shot, sometimes it’s because the ball is blocked by the opposing player. We modify “(block by SOMEONE)” to “blocked by SOMEONE”.
Events Foul Rebound Violation Timeout Freethrow Entergame Turnover Jumpball Shot
num. 986 2649 68 271 1103 1189 774 49 4400
Table 5: Statistics of the events in MbgKG. “num.” is the abbreviation of “number”. Numbers in bold indicate the highest two values.
Appendix C Evaluation Metric
Our task is oriented towards basketball live text broadcast, where concise and succinct text descriptions can allow users to quickly grasp and understand the content of the game in a short time. The names of the players and their corresponding fine-grained actions are the main components of the text descriptions. Existing evaluation metrics based on n-gram such as CIDEr, METEOR, Rouge-L, and BLEU, primarily penalize the semantically correct but textually inconsistent text descriptions. Each evaluation metric has its unique calculation method and principle, and no single metric can be considered as authoritative and comprehensive. Therefore, the most direct approach is to evaluate the quality of the text description by combining multiple metrics.
Since existing evaluation metrics cannot assess the accuracy of player name prediction, we introduce $e_{acc}$ to address this issue. When evaluating concise texts, the effectiveness of Rouge-L diminishes as it relies on the length of the longest common subsequence between the two texts. Therefore, compared to other evaluation metrics, the importance of Rouge-L would be somewhat lessened. We believe that the assessment of player name prediction is crucial, as it is a fundamental requirement in real-world applications. Based on the aforementioned analysis, we assign the highest coefficient to $e_{acc}$ and the lowest coefficient to Rouge-L. The coefficients of CIDEr and METEOR fall between those of $e_{acc}$ and Rouge-L. From experience, the coefficient of eacc is set to 4, the coefficients of CIDEr and METEOR are set to 2, and the coefficient of Rouge-L is set to 1. We sought opinions from several experts and scholars in the field of video captioning, all of whom expressed that the coefficient allocation has a certain degree of rationality.
Appendix D Qualitative Results
Fig. 10 shows the qualitative results on MbgVC dataset, including generation results of V2C, Clip4Caption, SwinBert and our EAC model. V2C, Clip4Caption and SwinBert can correctly generate some actions but fail to generate correct entity names because these names never appear during training. However, with the help of the external knowledge, our EAC could well relate entities in video clips with the fine-grained actions. These cases demonstrate that our newly proposed EAC model more consistently with the requirement of practical application.

### 28 Feb 2024 v2

@@
-Knowledge Graph Supported Benchmark and Video Captioning for Basketball
+Knowledge Guided Entity-aware Video Captioning and A Basketball Benchmark
 Abstract
 Despite the recent emergence of video captioning models, how to generate the text description with specific entity names and fine-grained actions is far from being solved, which however has great applications such as basketball live text broadcast.
-In this paper, a new multimodal knowledge supported basketball benchmark for video captioning is proposed.
-Specifically, we construct a Multimodal Basketball Game Knowledge Graph (MbgKG) to provide knowledge beyond videos.
-Then, a Multimodal Basketball Game Video Captioning (MbgVC) dataset that contains 9 types of fine-grained shooting events and 286 players’ knowledge (i.e., images and names) is constructed based on MbgKG.
-We develop a novel framework in the encoder-decoder form named Entity-Aware Captioner (EAC) for basketball live text broadcast.
-The temporal information in video is encoded by introducing the bi-directional GRU (Bi-GRU) module.
-And the multi-head self-attention module is utilized to model the relationships among the players and select the key players.
-Besides, we propose a new performance evaluation metric named Game Description Score (GDS), which measures not only the linguistic performance but also the accuracy of the names prediction.
-Extensive experiments on MbgVC dataset demonstrate that EAC effectively leverages external knowledge and outperforms advanced video captioning models.
-The proposed benchmark and corresponding codes will be publicly available soon.
-1 Introduction
+In this paper, a new multimodal knowledge graph supported basketball benchmark for video captioning is proposed.
+Specifically, we construct a multimodal basketball game knowledge graph (KG_NBA_2022) to provide additional knowledge beyond videos.
+Then, a multimodal basketball game video captioning (VC_NBA_2022) dataset that contains 9 types of fine-grained shooting events and 286 players’ knowledge (i.e., images and names) is constructed based on KG_NBA_2022.
+We develop a knowledge guided entity-aware video captioning network (KEANet) based on a candidate player list in encoder-decoder form for basketball live text broadcast.
+The temporal contextual information in video is encoded by introducing the bi-directional GRU (Bi-GRU) module.
+And the entity-aware module is designed to model the relationships among the players and highlight the key players.
+Extensive experiments on multiple sports benchmarks demonstrate that KEANet effectively leverages extera knowledge and outperforms advanced video captioning models.
+The proposed dataset and corresponding codes will be publicly available soon.
+Index Terms:
+Knowledge Graph, Video captioning, Entity-aware, Basketball.
+I Introduction
 Video captioning (VC) is a crucial computer vision task that requires the model to outputs corresponding text descriptions based on the given video.
-Despite the recent rapid development of video captioning [17, 12, 16, 13, 36, 33], how to automatically generate a text description with specific entity names and fine-grained actions is seldom investigated.
+By associating visual information and text elements, this task has been blossoming and gaining increasing attention owing to its many promising applications, such as video title automatic generation [1], visually-impaired assistance [2], video storytelling [3] and online video search [4, 5].
+Despite the recent rapid development of video captioning [6, 7, 8, 9, 10, 11, 12], these methods are not practical in real-world scenarios.
+They cannot automatically generate a text description with specific entity names and fine-grained actions.
 As the basketball live text broadcast example in Fig.
@@
 In contrast, if the model has game-related knowledge such as players who appear in the game and fine-grained actions in basketball, it can generate the knowledge-grounded text description (e.g., Brandon Ingram misses the 2pt jump shot and Justise Winslow gets the defensive rebound).
+In addition, existing common used benchmarks, including MSVD [13], YouCook [14], MSR-VTT [15], and ActivityNet Captions [16], simplify the task by using indefinite pronouns like “a man”, “a woman” or “a group of men” instead of specific entity names.
+And the actions in the annotated captions are coarse-grained.
+These benchmarks cannot provide relevant knowledge beyond videos and fine-grained action annotations to develop models to generate text descriptions with specific entity names and fine-grained actions.
 Figure 1: Comparison of conventional captioning with knowledge-grounded captioning.
@@
 And the fine-grained actions are marked green.
-In addition, existing common used benchmarks, including MSVD [2], YouCook [3], MSR-VTT [31], and ActivityNet Captions [9], simplify the task by using indefinite pronouns like “a man”, “a woman” or “a group of men” instead of specific entity names.
-And the actions in the annotated captions are coarse-grained.
-These benchmarks cannot provide relevant knowledge beyond videos and fine-grained action annotations to develop models to generate text descriptions with specific entity names and fine-grained actions.
-To our best knowledge, there is no relevant benchmark in basketball that can support this task.
+An increasing number of researchers have discovered that conventional methods and benchmarks for video captioning fail to meet the requirements of practical applications and have made numerous attempts to address this issue.
+Mkhallati et al.
+[17] publicly release SoccerNet-Caption, the first dataset for dense video captioning in soccer broadcast videos.
+Although this work focuses on researching entity-aware video captioning, it cannot provide any additional knowledge for the model to generate text descriptions with specific entity names.
+Ayyubi et al.
+[18] first train an entity perception detector to detect entities in video frames.
+They then utilize a large language model to retrieve relevant knowledge and enrich the visual content, thereby generating entity-aware text descriptions for news videos summary.
+However, this approach is limited by the performance of entity perception detector and large language model.
+Furthermore, Qi et al.
+[19] build a unimodal knowledge graph and a benchmark (Goal) for soccer commentary.
+The provided knowledge is not comprehensive and lacks visual aspects.
+It can be seen that these methods are tailored to different domains.
+The introduction of relevant background knowledge is essential to achieve text descriptions with specific entity names in a particular field.
 Figure 2: An example of a Multimodal Basketball Game Knowledge Graph.
-To narrow this gap, a new multimodal knowledge graph supported basketball benchmark for video captioning is proposed.
+In this work, we focus on the task of knowledge guided entity-aware basketball video captioning (KEBVC), which requires the model to comprehend the content of the provided video and generate text descriptions that include specific entity names and fine-grained actions based on the additional knowledge.
+We propose a new multimodal knowledge graph supported video captioning benchmark for basketball live text broadcast.
 This benchmark requires the model to comprehend the video’s content and generate event descriptions that incorporate specific entities with additional knowledge.
 Specifically, we collect 25 NBA full-games multimodal data from professional basketball platforms in 2022-2023 season, including events, players’ information, teams’ information, and videos.
-Based on the collected data, we construct a Multimodal Basketball Game Knowledge Graph (MbgKG) to provide the knowledge beyond videos, as depicted in Fig.
+Based on the collected data, we construct a multimodal basketball game knowledge graph named KG_NBA_2022 to provide the knowledge beyond videos, as depicted in Fig.
 2.
-Subsequently, a Multimodal Basketball Game Video Captioning (MbgVC) dataset is constructed based on nodes in MbgKG and relationships among the selected nodes.
-MbgVC dataset comprises 9 types of basketball shooting events and 286 players’ knowledge (i.e., images and names), with data samples illustrated in Fig.
+Subsequently, a multimodal basketball game video captioning dataset named VC_NBA_2022 is constructed based on nodes in KG_NBA_2022 and relationships among the selected nodes.
+VC_NBA_2022 dataset comprises 9 types of basketball shooting events and 286 players’ knowledge (i.e., images and names), with data samples illustrated in Fig.
 3.
-To effectively utilize knowledge in the MbgVC dataset, we propose an encoder-decoder framework named Entity-Aware Captioner (EAC) for basketball live text broadcast.
-EAC is comprised of 3 separate unimodal encoders for videos, players’ images, and players’ names, as well as a pre-trained language model that serves as the text decoder for generating descriptions of the given videos.
-Moreover, the Bi-GRU module is introduced to encode the temporal information, while the multi-head self-attention module is introduced to model the associations among the candidate players and select the key players.
-In addition, we propose a new performance evaluation metric, the Game Description Score (GDS), which takes into account both the linguistic quality of the description (as used in conventional video captioning tasks) and the accuracy of name predictions.
+To effectively utilize knowledge in the VC_NBA_2022 dataset, we propose a knowledge guided entity-aware network (KEANet) based on a candidate player list in encoder-decoder form for basketball live text broadcast.
+KEANet is comprised of 3 separate unimodal encoders for videos, players’ images, and players’ names, as well as a pre-trained language model that serves as the text decoder for generating descriptions of the given videos.
+Moreover, the Bi-GRU module is introduced to encode the temporal contextual information, while the entity-aware module is designed to model the associations among the candidate players and highlight the key players.
 The main contributions of this paper are as follows:
-  • A Multimodal Basketball Game Knowledge Graph (MbgKG) is constructed to provide additional contextual knowledge that extends beyond the video contents.
-  • Based on MbgKG, nodes and relationships are extracted to construct our Multimodal Basketball Game Video Captioning (MbgVC) benchmark.
-Additionally, we propose a new evaluation metric for MbgVC that offers a more accurate and comprehensive evaluation by incorporating the accuracy of name predictions.
-  • We develop an Entity-aware Video Captioning framework in encoder-decoder form, which selectly incorporates key information from external knowledge to generate text descriptions with specific entity names.
-  • On the MbgVC, our proposed model outperforms existing advanced models, achieving leading performance.
-Figure 3: Data samples from the proposed dataset.
+  • We provide an in-depth analysis to discover characteristics of basketball domain and construct a multimodal Basketball Game Knowledge Graph.
+This knowledge graph introduce additional knowledge that extends beyond the video contents.
+Based on this Knowledge Graph, nodes and relationships are extracted to construct our Multimodal Basketball Game Video Captioning benchmark.
+  • We develop a knowledge guided entity-aware video captioning network based on a candidate player list in encoder-decoder form, which attentively incorporates key information from additional knowledge to generate text descriptions with specific entity names.
+  • To validate the generalization of the proposed model, we conduct experiments on the proposed basketball dataset and the football domain dataset Goal [19].
+The proposed model outperforms existing advanced models, achieving leading performance.
+Figure 3: Data sample from the proposed dataset.
 Each video is annotated by fileid, action type, caption, player images and player names.
 Each of the players involved in caption as well as their teammates serve as candidate players.
-2 Related Works
-2.1 Computer Vision with Knowledge Graph
+II Related Works
+II-A Computer Vision with Knowledge Graph
 A knowledge graph (KG) is essentially a large-scale semantic network that contains entities, concepts as nodes and various semantic relationships among them as edges [20].
@@
 One of the main challenges in video-text retrieval task is to identify fine-grained semantic associations between video and text.
-To address this issue, under the guidance of external knowledge, Fang et al.
+To address this issue, under the guidance of additional knowledge, Fang et al.
 [22] utilize the associations between concepts to extend new concepts and enrich the representation of videos.
@@
 Gu et al.
-[6] propose a text with knowledge graph augmented transformer, which integrates external knowledge and leverages multimodal information to mitigate the challenges posed by long-tail words.
+[23] propose a text with knowledge graph augmented transformer, which integrates additional knowledge and leverages multimodal information to mitigate the challenges posed by long-tail words.
 However, unlike the aforementioned works that extract knowledge from unimodal knowledge graphs, our approach extracts knowledge from a multimodal knowledge graph, utilizing this multimodal knowledge to assist the model in generating descriptions that include specific entity names and fine-grained actions.
-2.2 Video Captioning Benchmarks
+II-B Video Captioning
+Video captioning is a crucial task in video understanding, where models generate text descriptions for given videos.
+Some works [24, 25, 8, 26, 27, 28, 29, 30] utilize a visual encoder to extract video representations from a set of video frames, and a language decoder transfer these representations to the corresponding descriptions.
+Recently, CLIP (Contrastive Language-Image Pretraining) [31] has demonstrated its superior performance on various visual-linguistic tasks relying on large-scale contrastive pre-training with image-text pairs.
+Therefore, Clip4Caption [32] employs CLIP to acquire the aligned visual-text representation, leading to significant improvements in video captioning performance.
+Benefiting from the flexibility of the transformer [33] architecture, SwinBert [7] introduces Video Swin Transformer [34] as the video encoder to encode spatial-temporal representations from video frames.
+None of the previously mentioned works can address the challenge of generating action entity names effectively.
+In stark contrast, we propose to incorporate additional knowledge based on an encoder-decoder model structure to generate text descriptions with specific entity names and fine-grained actions for basketball live text broadcasts, demonstrating the practicality of our entity-aware video captioning in real applications.
+II-C Video Captioning Benchmarks
 Existing widely used benchmarks to support the video captioning task include MSVD [13], MSR-VTT [15], YouTube Hightlight [35] and VATEX [36].
@@
 [39] construct a large-scale Chinese movie benchmark, which requires models to generate role-aware narration paragraphs when there are no actors speaking.
+Byeong et al.
+[40] propose a benchmark for automatically generating commentary on baseball games.
+The descriptions in this benchmark focus on player categories rather than specific player names.
+To achieve meaningful news summarization, Ayyubi et al.
+[18] propose the task of summarizing news video directly to entity-aware captions.
+They also release a large-scale dataset, to support research on this task.
+Mkhallati et al.
+[17] release a benchmark for soccer broadcasts commentaries.
+Although the text caption in the benchmark contains specific player names, they do not provide any experimental support for generating a caption with specific names.
+On the contrary, each specific name is replaced by a specific token (e.g., [TEAM], [COACH], [REFEREE], and [PLAYER]).
+Figure 4: Illustration of extracting relevant data using relationship extraction from the knowledge graph and constructing the dataset.
 Qi et al.
@@
 This indiscriminate inclusion could introduce noise, affecting caption accuracy.
-Contrastingly, our approach first employs a multimodal knowledge graph, integrating both textual and visual information, such as player names and images.
+Contrastingly, our approach first employs a multimodal knowledge graph, integrating both text and visual information, such as player names and images.
 This enriches the context for video captioning, improving the match with video content.
-Second, we propose a selective attention mechanism, modeling associations among candidate players and incorporating key players’ knowledge.
+Second, we propose a entity-aware module, modeling associations among candidate players and incorporating key players’ knowledge.
 This effectively reduces noise and enhances caption precision.
-These advancements distinguish our work from Qi et al.’s [19] and other existing models, demonstrating the potential of multimodal knowledge and selective attention in video captioning tasks.
-2.3 Video Captioning Models
-Figure 4: Illustration of the proposed video captioning model that leverages external knowledge from multimodal knowledge graph.
-Video captioning is a crucial task in video understanding, where models generate text descriptions for given videos.
-Some works [10, 14, 16, 36] utilize a visual encoder to extract video representations from a set of video frames, and a language decoder transfer these representations to the corresponding descriptions.
-Recently, CLIP (Contrastive Language-Image Pretraining) [20] has demonstrated its superior performance on various visual-linguistic tasks relying on large-scale contrastive pre-training with image-text pairs.
-Therefore, Clip4Caption [26] employs CLIP to acquire the aligned visual-text representation, leading to significant improvements in video captioning performance.
-Benefiting from the flexibility of the transformer [27] architecture, SwinBert [12] introduces Video Swin Transformer [15] as the video encoder to encode spatial-temporal representations from video frames.
-However, none of the previously mentioned works can address the challenge of generating action entity names effectively.
-In stark contrast, we propose to incorporate external knowledge based on an encoder-decoder model structure to generate text descriptions with specific entity names and fine-grained actions for basketball live text broadcasts, demonstrating the practicality of our entity-aware video captioning in real applications.
-3 Proposed Benchmark
+These advancements distinguish our work from Qi et al.’s [19] and other existing models, demonstrating the potential of multimodal knowledge and entity-aware in video captioning tasks.
+III Proposed Benchmark
 In this section, we first construct a multimodal knowledge graph as the knowledge base, and then extract relevant data and knowledge to construct a multimodal basketball dataset.
-Furthermore, a new metric for this dataset is proposed to assess the performance of models.
-3.1 Multimodal Basketball Game Knowledge Graph
+III-A Multimodal Basketball Game Knowledge Graph
 Our knowledge graph is based on basketball game videos with corresponding event descriptions.
@@
 (2) Parse and structure the descriptions of each type of event.
-(3) Match these event descriptions with the corresponding video timestamps using OCR toolkits.
+(3) Match these event descriptions with the corresponding video timestamps.
+By analyzing the play-by-play data, game events can be divided into 9 categories, including “Foul”, “Rebound”, “Violation”, “Timeout”, “Freethrow”, “Enter game”, “Turnover”, “Jump ball”, and “Shot”.
+Each text description has its owns keyword, such as the description of the “Foul” event “Personal foul by G.
+Temple (drawn by A.
+Drummond)”, where the keyword is “foul”.
+Therefore, the type of this event is “Foul”.
+For this semi-structured data, the sentence is parsed into multi-tuples using the character index of the keyword and specific words that appear in the sentence, such as “drawn by”.
+To incorporate video information into the multi-tuples, the video timestamp is intended to be associated with the text description.
+OCR is used to identify the time on the scoreboard in each frame and record the timestamp of the current frame.
+Specifically, we employ the open-source OCR toolkit PaddleOCR [^3] https://github.com/PaddlePaddle/PaddleOCR and Tesseract-OCR [41] to recognize the time simultaneously.
+By using multiple OCR toolkits, we can improve the overall accuracy of recognition.
+This is because one toolkit may succeed where the other fails.
+In addition to event information, player and team information are also stored in multi-tuples.
+Through these efforts, a Multimodal Basketball Game Knowledge Graph containing 11489 events and 42870 relationships is constructed.
+Fig .2 shows an example of the Multimodal Basketball Game Knowledge Graph (MbgKG).
+TABLE I: Statistics of the events in KG_NBA_2022. “num.” is the abbreviation of “number”. Numbers in bold indicate the highest two values.
+Events Foul Rebound Violation Timeout Freethrow Entergame Turnover Jumpball Shot
+num. 986 2649 68 271 1103 1189 774 49 4400
+III-B Multimodal Basketball Video Game Captioning Dataset
+To construct a dataset that contains videos, text descriptions, and additional knowledge (i.e., player images and names), the pertinent data is extracted from KG_NBA_2022 by utilizing the relations within KG_NBA_2022.
+In basketball, shot and rebound events are the most prevalent.
+Meanwhile, we count the number of different types of events in KG_NBA_2022.
+Table I shows that there are 4400 shot events and 2649 rebound events, which are the most common types of events.
+Based on the aforementioned common knowledge and statistical data, videos and event descriptions about shot and rebound are extracted from KG_NBA_2022 to construct the dataset.
+To be more realistic, the text descriptions need to be further modified.
+Figure 5: Word cloud of VC_NBA_2022 and Goal datasets. The bigger the font, the more percentage it occupies.
+When a shot fails to score in basketball, it is often followed by a rebound event.
+In line with this pattern, we have combined the shot and rebound into a single event.
+Shots in basketball can be categorized as two-point (2pt) shots, three-point (3pt) shots, and layups.
+Rebounds are further divided into defensive (def.) rebounds and offensive (off.) rebounds.
+The classification scheme of our dataset is based on the NBA dataset [42] used for group activity recognition.
+However, unlike the NBA dataset, where all videos are 6 seconds long and divided into 72 frames, we consider that different events may have different durations.
+Consequently, in our dataset, videos of different lengths are uniformly divided into 72 frames.
+Additionally, we have incorporated text descriptions and player information into our dataset.
+As shown in Fig.
+4, the event-action relation in the graph is used to extract 9 kinds of events, and the video-time relation in the graph is used to obtain the video timestamp of the event.
+Specifically, we first roughly estimate the start and end timestamps of the event based on the existing timestamp.
+Each clip undergoes a manual review process to ensure accurate start and end timestamps.
+Subsequently, the video-description relation is used to obtain the text description, which is then matched with the corresponding video clip.
+We extract player-related information (player images and player names) from KG_NBA_2022 through the team-player and player-name relations.
+Since each event involves certain players, only the involved individuals and their teammates need to be considered as candidate players for video annotation.
+Using the sentence “B.
+Ingram misses 2-pt jump shot from 19 ft and defensive rebound by J.
+Winslow” as an example, we will explain our text modification process.
+Due to the difficulty in generating distance in video captioning task, the “from 19 ft” in the sentence is removed.
+Based on the attribute information of the player in KG_NBA_2022, the abbreviated name (“J.
+Winslow”) in the sentence is replaced with the full name (“Justise Winslow”).
+This change might increase the complexity of name generation but is more aligned with real-world scenarios.
+To improve text fluency, alterations are made based on the character index.
+Specifically, we identify the index of keywords like ”defensive” and apply a rule (“defensive/offensive rebound by SOMEONE” $\Longrightarrow$ “SOMEONE gets the defensive/offensive rebound”) to revise the sentence.
+Finally, there are a total 3977 videos [^4] The scale of the current dataset is not large enough, which is emergent to be enriched in our subsequent research., each of which belongs to one of 9 kinds of events.
+Each video has one text description and several candidate players information (images and names).
+We randomly select 3162 clips for training and 786 clips for testing.
+We also provide word-cloud-based statistics in Fig.
+5 (a) to reveal the relative amount of different words.
+It shows that the top-4 subjects in VC_NBA_2022 are “jump”, “shot”, “3pt”, and “defensive”, followed by “rebound”, “2pt”, “layup”, and “makes”.
+Table II shows the sample distributions across different labels of events.
+In Table III, the comparison of our dataset with MSR-VTT, YouCook and ActivityNet Captions further demonstrates the fine-grained details of our captioning annotations.
+VC_NBA_2022 has the most sentences per second of 0.182, while the other datasets are all below 0.1, this reflects that our dataset has more detailed information in descriptions.
+Moreover, VC_NBA_2022 has 1.73 verbs in a sentence on average, higher than 1.41 for ActivityNet Captioning and 1.37 for MSR-VTT.
+Similarly, the verb ratio of VC_NBA_2022 which is computed by dividing the total number of verbs by the total number of words in the sentence is also much higher than other 3 datasets.
+This highlights that our dataset is primarily focused on the fine-grained actions of the subjects, aligning with our original intent.
+Figure 6: The architecture of knowledge guided entity-aware video captioning network (KEANet) based on a candidate player list in encoder-decoder form.
+TABLE II: Statistics of the labels in VC_NBA_2022. “succ.” and “fail.” are abbreviations of “success” and “failure”, respectively.
 Labels
@@
 Test num. 95 41 95 108 32 67 125 61 162
-Table 1: Statistics of the labels in NBA. “succ.” and “fail.” are abbreviations of “success” and “failure”, respectively.
-By analyzing the play-by-play data, game events can be divided into 9 categories, including “Foul”, “Rebound”, “Violation”, “Timeout”, “Freethrow”, “Enter game”, “Turnover”, “Jump ball”, and “Shot”.
-Each text description has its owns keyword, such as the description of the “Foul” event “Personal foul by G.
-Temple (drawn by A.
-Drummond)”, where the keyword is “foul”.
-Therefore, the type of this event is “Foul”.
-For this semi-structured data, the sentence is parsed into multi-tuples using the character index of the keyword and specific words that appear in the sentence, such as “drawn by”.
-To incorporate video information into the multi-tuples, the video timestamp is intended to be associated with the text description.
-OCR is used to identify the time on the scoreboard in each frame and record the timestamp of the current frame.
-Specifically, we employ the open-source OCR toolkit PaddleOCR [^3] https://github.com/PaddlePaddle/PaddleOCR and Tesseract-OCR [24] to recognize the time simultaneously.
-By using multiple OCR toolkits, we can improve the overall accuracy of recognition.
-This is because one toolkit may succeed where the other fails.
-In addition to event information, player and team information are also stored in multi-tuples.
-Through these efforts, a Multimodal Basketball Game Knowledge Graph containing 11489 events and 42870 relationships is constructed.
-An example of the Multimodal Basketball Game Knowledge Graph (MbgKG) is shown in Fig.
-2.
-More details on the construction of MbgKG can be found in Appendix A.
-3.2 Multimodal Basketball Video Game Captioning Dataset
-To construct a dataset that contains videos, text descriptions, and external knowledge (i.e., player images and names), the pertinent data is extracted from MbgKG by utilizing the relations within MbgKG.
-As shot and rebound events are the most prevalent in basketball, we only extract the data about these events from MbgKG.
-To be more realistic, the text descriptions need to be further modified.
-When a shot fails to score in basketball, it is often followed by a rebound event.
-In line with this pattern, we have combined the shot and rebound into a single event.
-Shots in basketball can be categorized as two-point (2pt) shots, three-point (3pt) shots, and layups.
-Rebounds are further divided into defensive (def.) rebounds and offensive (off.) rebounds.
-The classification scheme of our dataset is based on the NBA dataset [32] used for group activity recognition.
-However, unlike the NBA dataset, where all videos are 6 seconds long and divided into 72 frames, we consider that different events may have different durations.
-Consequently, in our dataset, videos of different lengths are uniformly divided into 72 frames.
-Additionally, we have incorporated text descriptions and player information into our dataset.
-As shown in Fig.
-4, the event-action relation in the graph is used to extract 9 kinds of events, and the video-time relation in the graph is used to obtain the video timestamp of the event.
-Specifically, we first roughly estimate the start and end timestamps of the event based on the existing timestamp.
-Each clip undergoes a manual review process to ensure accurate start and end timestamps.
-Subsequently, the video-description relation is used to obtain the text description, which is then matched with the corresponding video clip.
-We extract player-related information (player images and player names) from MbgKG through the team-player and player-name relations.
-Since each event involves certain players, only the involved individuals and their teammates need to be considered as candidate players for video annotation.
-Using the sentence “B.
-Ingram misses 2-pt jump shot from 19 ft and defensive rebound by J.
-Winslow” as an example, we will explain our text modification process.
-Due to the difficulty in generating distance in video captioning task, the “from 19 ft” in the sentence is removed.
-Based on the attribute information of the player in MbgKG, the abbreviated name (“J.
-Winslow”) in the sentence is replaced with the full name (“Justise Winslow”).
-This change might increase the complexity of name generation but is more aligned with real-world scenarios.
-To improve text fluency, alterations are made based on the character index.
-Specifically, we identify the index of keywords like ”defensive” and apply a rule (“defensive/offensive rebound by SOMEONE” $\Longrightarrow$ “SOMEONE gets the defensive/offensive rebound”) to revise the sentence.
-Finally, there are a total 3977 videos [^4] The scale of the current dataset is not large enough, which is emergent to be enriched in our subsequent research., each of which belongs to one of 9 kinds of events.
-Each video has one text description and several candidate players information (images and names).
-We randomly select 3162 clips for training and 786 clips for testing.
-Tab.
-1 shows the sample distributions across different labels of events.
-More details on the construction of MbgVC can be found in Appendix B.
-In Tab.
-2, the comparison of our dataset with MSR-VTT, YouCook and ActivityNet Captions further demonstrates the fine-grained details of our captioning annotations.
-MbgVC has the most sentences per second of 0.182, while the other datasets are all below 0.1, this reflects that our dataset has more detailed information in descriptions.
-Moreover, MbgVC has 1.73 verbs in a sentence on average, higher than 1.41 for ActivityNet Captioning and 1.37 for MSR-VTT.
-Similarly, the verb ratio of MbgVC which is computed by dividing the total number of verbs by the total number of words in the sentence is also much higher than other 3 datasets.
-This highlights that our dataset is primarily focused on the fine-grained actions of the subjects, aligning with our original intent.
-3.3 Evaluation Metric
-Existing video captioning benchmarks mainly adopt n-gram based metrics, including CIDEr [28], METEOR [1], Rouge-L [11], and BLEU [18].
-BLEU evaluates the quality of the generated text by calculating the n-gram overlap between the generated text and the reference text.
-ROUGE-L pays more attention to generating the longest common subsequence between the generated text and the reference text, and evaluates sentence quality in this way.
-METEOR takes more factors into account, including stem matching, synonym matching, and word order.
-CIDEr uses TF-IDF [22] to assign different weights to n-grams of different lengths, then calculates the cosine similarities of n-grams between generated texts and reference texts and averages them to get the final score.
-These metrics primarily emphasize the consistency of text rather than the accuracy of semantics.
-In other words, if the generated text differs from the reference in structure or vocabulary, it may be penalized by these n-gram based evaluation metrics, even if the conveyed information is completely correct.
-However, for entity names, they cannot be replaced with the similar words.
-To take into account the predictive accuracy of specific entity names, the entity name prediction accuracy $e_{acc}$ is introduced to accurately evaluate the accuracy of name prediction.
-$e_{acc}$ is a score calculated by dividing the number of correctly predicted names by the total number of names.
-Finally, we settle on a new metric, Game Description Score (GDS), as follows:
-$gds=\frac{4\cdot e_{acc}+2\cdot cs+2\cdot ms+1\cdot rs}{9}\times 100,$ (1)
-where $cs$, $ms$, and $rs$ refer to entity name prediction accuracy score, CIDEr score, METEOR score, and Rouge-L score, respectively.
-Since CIDEr involves the partitioning and matching of n-grams, BLEU is not included in the calculation.
-More detailed descriptions of GDS can be found in Appendix C.
+TABLE III: Comparisons of different video caption datasets. Numbers in bold indicate the highest value.
 Dataset
@@
 0.028 1.41 10.4%
-MbgVC 0.182 1.73 14.8%
-Table 2: Comparisons of different video caption datasets. Numbers in bold indicate the highest value.
-4 Knowledge Graph Supported Basketball Video Captioning
-4.1 Problem Formulation
-To generate precise and concise textual descriptions of the basketball live text broadcast, we introduce the concept of a Knowledge-Aided Basketball Video Captioning (KABVC) task.
-In this task, the model is required to comprehend the content of the provided video and generate text descriptions that include specific entity names and fine-grained actions based on the external knowledge.
+VC_NBA_2022 0.182 1.73 14.8%
+IV Knowledge-based Entity-aware Basketball Video Captioning
+IV-A Problem Formulation
+To generate precise and concise text descriptions of the basketball live text broadcast, we introduce the concept of KEBVC task.
+In this task, the model is required to comprehend the content of the provided video and generate text descriptions that include specific entity names and fine-grained actions based on the additional knowledge.
 This task can be formulated as: given the basketball video $V_{b}$, the objective is to select video-related player knowledge $K_{p}=s\left(V_{b}\right)$ and generate the text description $D_{v,k}$.
-${\rm KABVC}:D_{v,k}=m\left(V_{b},K_{p}\right)=m\left(V_{b},s\left(V_{b}\right)\right),$ (2)
+${\rm KEBVC}:D_{v,k}=m\left(V_{b},K_{p}\right)=m\left(V_{b},s\left(V_{b}\right)\right),$ (1)
 where $s\left(\cdot\right)$ and $m\left(\cdot\right)$ denote the models’ abilities on aligning video and knowledge (player information) and transferring video and knowledge to the text description, respectively.
-Model CIDEr METEOR Rouge-L BLEU-1 BLEU-2 BLEU-3 BLEU-4 GDS
-V2C 13.7 18.7 45.9 50.1 39.5 27.0 14.9 12.3
-Clip4Caption 19.4 26.7 51.2 49.1 42.5 35.4 28.8 25.0
-SwinBert 20.1 26.5 49.0 47.8 41.4 34.5 28.4 25.6
-EAC 138.5 28.0 54.9 53.1 46.4 38.8 32.4 56.9
-Table 3: Combined inference performance on MbgVC. Numbers in bold indicate the best performance.
-4.2 Proposed Model
-For the ${\rm KABVC}$ task, we propose a model with an encoder-decoder framework named Entity-Aware Captioner (EAC).
-The overall structure of EAC framework is shown in Fig.
-4.
-Given the raw video frames which are of size $T\times 3\times H\times W$, consisting of $T$ frames and each has $3\times H\times W$ pixels.
-We feed them into Resnet-18 [7], which is the video encoder in EAC, and extract the video feature $F_{T}\in{\rm R}^{T\times D_{r}\times h\times w}$.
-$D_{r}$ is the hidden size of Resnet-18.
-The global representations of $F_{T}$ are then fed into Bi-GRU module to further encoder temporal information and obtain the feature $F_{v}$.
+TABLE IV: Combined inference performance on VC_NBA_2022. Numbers in bold indicate the best performance.
+Model CIDEr METEOR Rouge-L BLEU-1 BLEU-2 BLEU-3 BLEU-4 RoleF1
+V2C 13.7 18.7 45.9 50.1 39.5 27.0 14.9 0.0
+Clip4Caption 70.4 26.7 51.2 49.1 42.5 35.4 28.8 8.3
+SwinBert 69.1 26.5 49.0 47.8 41.4 34.5 28.4 7.4
+KEANet 138.5 28.0 54.9 53.1 46.4 38.8 32.4 20.6
+Figure 7: Detailed architecture of the entity-aware module.
+IV-B Proposed Model
+For the ${\rm KEBVC}$ task, we propose a knowledge guided entity-aware video captioning network (KEANet) based on a candidate player list in encoder-decoder form.
+The overall structure of KEANet framework is shown in Fig.
+6.
+Given the raw video frames which are of size $T\times 3\times H\times W$, consisting of $T$ frames and each has $3\times H\times W$ pixels, we feed them into the CNN-based visual encoder in KEANet and extract the video feature $F_{T}\in{\rm\mathbb{R}}^{T\times D_{r}\times h\times w}$.
+$D_{r}$ is the hidden size of visual feature.
+The global representations of $F_{T}$ are then fed into Bi-GRU module to further encode temporal contextual information and obtain the feature $F_{v}\in{\rm\mathbb{R}}^{T\times D}$.
 $F_{v}=W_{1}\left({{\rm GRU}\left({\rm AvgPool}\left(F_{T}\right)\right)}\right),$ (2)
-where $W_{1}\in{\rm R}^{D_{r}\times D}$ is the linear mapping layer, which maps the video feature to text space.
+where $W_{1}\in{\rm\mathbb{R}}^{D_{r}\times D}$ is the linear mapping layer, which maps the video feature to text space.
 $D$ is the hidden size of the decoder module.
 ${\rm GRU}\left(\cdot\right)$ denotes the Bi-GRU module and ${\rm AvgPool}\left(\cdot\right)$ denotes the average pooling layer.
-Each video has the corresponding candidate $N$ players’ information as the external knowledge to assist the generation of text with the specific entity names.
+Each video has the corresponding candidate $N$ players’ information as the additional knowledge to assist the generation of text with the specific entity names.
 Each player image is of size $3\times H_{p}\times W_{p}$.
-For candidate players’ images, Resnet-18 is used to extract the features $F_{N}\in{\rm R}^{N\times D_{r}\times h_{p}\times w_{p}}$.
-The global features $F_{p}\in{\rm R}^{N\times D_{r}}$ of $N$ images are obtained by Eq.
-4.
+For candidate players’ images, we use the CNN-based visual encoder to extract the features $F_{N}\in{\rm\mathbb{R}}^{N\times D_{r}\times h_{p}\times w_{p}}$.
+The global features $F_{p}\in{\rm\mathbb{R}}^{N\times D}$ of $N$ images are obtained by (3).
 $F_{p}=W_{2}\left({\rm AvgPool}\left(F_{N}\right)\right)),$ (3)
-where $W_{2}\in{\rm R}^{D_{r}\times D}$ is the linear mapping layer, which maps the image feature to text space.
-The text encoder of large language model T5 [21] is used to transform candidate names into a sequence of embeddings $T_{n}\in{\rm R}^{N\times D}$.
-After that, the image features $F_{p}$ and name features $T_{n}$ are added by corresponding positions to obtain multimodal player features $F_{m}\in{\rm R}^{N\times D}$.
+where $W_{2}\in{\mathbb{R}}^{D_{r}\times D}$ is the linear mapping layer, which maps the image feature to text space.
+TABLE V: Combined inference performance on Goal. Numbers in bold indicate the best performance.
+Method CIDEr METEOR Rouge-L BLEU-1
+V2C 0.13 2.11 3.42 3.38
+Clip4Caption 2.17 5.01 5.45 5.71
+SwinBERT 2.20 5.12 5.32 5.65
+KEANet 3.68 6.38 10.45 14.88
+The text encoder of large language model T5 [43] is used to transform candidate names into a sequence of embeddings $T_{n}\in{\mathbb{R}}^{N\times D}$.
+After that, the image features $F_{p}$ and name features $T_{n}$ are added by corresponding positions to obtain multimodal player features $F_{m}\in{\mathbb{R}}^{N\times D}$.
 Basketball events involve interactions between multiple players.
 For example, in “Myles Turner makes the 2pt layup with an assist from Tyrese Haliburton”, “Myles Turner” scores a two-point layup with the assist of “Tyrese Haliburton”.
-Therefore, the multimodal player features $F_{m}$ are fed into the multi-head self-attention module to model the relationship among candidate players.
-The player features based on attention and the video features containing motion information are then concatenated and fed into the decoder ${\rm\Psi_{T5}}\left(\cdot\right)$ of T5 to obtain the text description $C$.
-$C={\rm\Psi_{T5}}\left(W_{3}\left({\rm Concat}\left[{\rm H_{M-ATT}}\left(F_{m}\right),F_{v}\right]\right)\right),$ (5)
-where $W_{3}\in{\rm R}^{D\times D}$ is the linear mapping layer, which maps the concatenated feature to the vector space of the T5 model.
-${\rm Concat}\left[,\right]$ denotes the Concatenation function in Python, and ${\rm H_{M-ATT}\left(\cdot\right)}$ denotes the multi-head self-attention module.
-The Bi-GRU module, 3 linear layers, multi-head self-attention and text decoder are trained to maximize the log-likelihood over the training set given by Eq. 6.
-$L_{\Theta}=\sum_{t=1}^{N_{C}}log{\rm P_{r}}\left(y_{t}|y_{t-1},\left[F_{m},F_{v}\right];\Theta\right),$ (6)
+Therefore, we design the entity-aware module to highlight key players from a candidate player list.
+The architecture of entity-aware module is shown in Fig.
+7.
+entity-aware module first uses the cross-attention to fuse player features and video features to obtain $F_{f}$.
+This connects the players to the video content, putting the players in a specific scene.
+$F_{a}$ is sent to the multi-head self-attention sub-module to model the relationship among candidate players.
+The player features based on attention and the video feature are then concatenated and fed into the decoder ${\rm\Psi_{T5}}\left(\cdot\right)$ of T5 to obtain the text description $C$.
+$C={\rm\Psi_{T5}}\left(W_{3}\left({\rm Concat}\left[{\rm E_{a}}\left(F_{m},F_{v}\right),F_{v}\right]\right)\right),$ (4)
+where $W_{3}\in{\mathbb{R}}^{D\times D}$ is the linear mapping layer, which maps the concatenated feature to the vector space of the T5 model.
+${\rm Concat}\left[,\right]$ denotes the concatenation function in Python, and ${\rm E_{a}\left(\cdot\right)}$ denotes the entity-aware module.
+The Bi-GRU module, 3 linear layers, entity-aware module and text decoder are trained to maximize the log-likelihood over the training set given by (5).
+$\mathcal{L}_{\Theta}=\sum_{t=1}^{N_{C}}log{\rm P_{r}}\left(y_{t}|y_{t-1},\left[F_{m},F_{v}\right];\Theta\right),$ (5)
 where $y_{t}$ denotes the one-hot vector probability of each word at time $t$. $N_{C}$ denotes the length of the caption. And $\Theta$ denotes the learnable parameters.
-5 Experiments
-To verify the performance of the proposed model, EAC is compared with the advanced video captioning models.
-We further conduct the ablation experiments to verify the effectiveness of each conponent in EAC.
-5.1 Implementation Details
-The proposed EAC uses ResNet-18 pretrained on the ImageNet dataset [23] as the visual encoder, which hidden size $D_{r}$ is 512.
-EAC is trained on the proposed MbgVC dataset with 100 training epochs.
-In MbgVC dataset, $T$ frames are sampled by using segment-based method [29].
-Each frame size is $720\times 1280$.
+V Experiments
+To verify the performance of the proposed model, KEANet is compared with the advanced video captioning models on VC_NBA_2022 and Goal [19].
+We further conduct the ablation experiments to verify the effectiveness of each conponent in KEANet.
+V-A Implementation Details
+The proposed KEANet uses ResNet-18 [44] pretrained on the ImageNet dataset [45] as the visual encoder, which hidden size $D_{r}$ is 512.
+KEANet is trained on the proposed VC_NBA_2022 and Goal datasets with 100 training epochs.
+In these datasets, $T$ frames are sampled by using segment-based method [46].
+Each frame size is $1280\times 720$.
+A high resolution of video frame is helpful for the model to understand the players’ action in the video.
 Concretely, $T$ is set to 18.
-The size of player’s image is $120\times 180$ and the number of candidate players in each video is not fixed.
+The size of player’s image is $180\times 120$ and the number of candidate players in each video is not fixed.
 For the large language model T5, the hidden size is 768.
@@
 Beam search with a beam size of 2 is used for inference.
-The proposed EAC is implemented with Python 3.9 and PyTorch 1.12, and is performed on a server with an Nvidia 3090ti GPU.
-5.2 Main Results
-EAC is compared with three advanced video captioning models, including V2C [5], Clip4Caption [26], and SwinBert [12].
+The proposed KEANet is implemented with Python 3.9 and PyTorch 1.12, and is performed on a server with an Nvidia 3090ti GPU.
+Figure 8: Qualitative results on VC_NBA_2022 dataset.
+(V2C: Video2Commonsense; C4C: Clip4Caption; SwB: SwinBert; KEA: our proposed model; GT: the ground truth).
+The different specific entity names are marked in red and blue, respectively.
+And the fine-grained actions are marked green.
+Since V2C does not have its own tokenizer and vocabulary list, it cannot decode names.
+So names are replaced with the special token $<$UNK$>$.
+V-B Evaluation metrics
+Existing video captioning benchmarks mainly adopt ngram-based metrics, including CIDEr [48], METEOR [49], Rouge-L [50], and BLEU [51].
+BLEU evaluates the quality of the generated text by calculating the n-gram overlap between the generated text and the reference text.
+Rouge-L pays more attention to generating the longest common subsequence between the generated text and the reference text, and evaluates sentence quality in this way.
+METEOR takes more factors into account, including stem matching, synonym matching, and word order.
+CIDEr uses TF-IDF [52] to assign different weights to n-grams of different lengths, then calculates the cosine similarities of n-grams between generated texts and reference texts and averages them to get the final score.
+These metrics primarily emphasize the consistency of text rather than the accuracy of semantics.
+In other words, if the generated text differs from the reference in structure or vocabulary, it may be penalized by these n-gram based evaluation metrics, even if the conveyed information is completely correct.
+However, for entity names, they cannot be replaced with the similar words.
+To accurately evaluate the performance of specific entity name generation, we introduce RoleF1 score (F1 score of entity name generation) from the work [39].
+For RoleF1, we extract entity names from the ground truth and the generated texts.
+Recall is used to measure how the generated texts covers the entities appearing in the game video clip.
+Taking into account that these generated entity names could also come from the model’s hallucinations, such as from incorrect games, we will also consider Precision.
+Finally, RoleF1 score is calculated with Precision and Recall.
+Figure 9: The relationship between metric CIDEr and RoleF1. C4C denotes Clip4Caption.
+V-C Results on VC_NBA_2022
+KEANet is compared with three advanced video captioning models, including V2C [37], Clip4Caption [32], and SwinBert [7].
 V2C is the Transformer-based model that generates relevant commonsense descriptions of the given video.
@@
 SwinBert introduces Video Swin Transformer to encode spatial-temporal representations from video frames.
-As shown in Tab.
-3, EAC outperforms other 3 models by a large margin on CIDEr.
+V2C, Clip4Caption, and SwinBert take only videos as input, and are all trained and tested on our proposed dataset.
+As shown in Table IV, KEANet outperforms other 3 models by a large margin on CIDEr.
 This notable performance can be attributed to the way CIDEr calculates cosine similarities of n-grams between the generated text and reference text.
@@
 V2C, Clip4Caption, and SwinBert exhibit lower performance in accurately generating names, resulting in much lower CIDEr scores.
-Although EAC achieves the best performance across all metrics, the differences are not as pronounced in METEOR, Rouge-L, and BLEU.
+Although KEANet achieves the best performance across all metrics, the differences are not as pronounced in METEOR, Rouge-L, and BLEU.
 This can be attributed to the concise nature of the text descriptions in the dataset, primarily consisting of names and fine-grained actions.
-Notably, EAC outperforms SwinBert 31.3% on GDS and outperforms Clip4Caption 31.9% on GDS.
-The higher GDS indicates better overall model performance, particularly in name prediction.
+Notably, KEANet outperforms SwinBert 13.2% on RoleF1 and outperforms Clip4Caption 12.3% on RoleF1.
+The higher RoleF1 indicates better name prediction.
 These results underscore the capability of our model to generate accurate names and fine-grained actions in live text broadcast task.
-Qualitative results can be found in Appendix D.
-Model Ki Kn Attn. Bi-GRU CIDEr METEOR Rouge-L BLEU-4 GDS
-20.1 20.6 40.9 23.0 19.1
-18.0 20.0 38.0 23.2 16.3
-110.6 25.2 49.1 27.3 46.7
-115.7 25.6 53.1 28.1 49.8
-119.3 26.8 53.3 29.5 51.4
-122.2 27.3 54.6 30.5 52.5
-138.5 28.0 54.9 32.4 56.9
-Table 4: Ablation study on MbgVC.
+Fig.
+8 shows the qualitative results on VC_NBA_2022 dataset, including generation results of V2C, Clip4Caption, SwinBert and our KEANet model.
+V2C, Clip4Caption and SwinBert can correctly generate some actions but fail to generate correct entity names because these names never appear during training.
+However, with the help of the additional knowledge, our KEANet could well relate entities in video clips with the fine-grained actions.
+These cases demonstrate that our newly proposed KEANet model more consistently with the requirement of practical application.
+Further, we study the relation between the metric CIDEr and RoleF1.
+To this end, we plot the performance of KEANet and Clip4Caption with different RoleF1 scores in Fig.
+9.
+The plot shows that this relationship is approximately linear: more accurate entity name generation imply better model performance.
+V-D Results on Goal
+Goal is a benchmark which contains over 8.9k soccer video clips, 22k sentences and 42k knowledge triples.
+These sentences were converted from the commentator’s audios.
+This type of comment sentences have a certain degree of colloquialism and are relatively long.
+Therefore, this dataset poses certain challenges.
+On this basis, we filter out sentences and videos that do not contain entity names.
+We modify the format of the dataset to be the same as VC_NBA_2022.
+In addition, we save the names of teams and players in the dataset.
+The revised Goal dataset is shown in Fig.
+10.
+KEANet is compared with three advanced video captioning models, including V2C, Clip4Caption, and SwinBert.
+V2C, Clip4Caption, and SwinBert take only videos as input, and are all trained and tested on the revised Goal dataset.
+As shown in Table V, we compare the performance of V2C, Clip4Caption, SwinBERT, and KEANet on several metrics including CIDEr, METEOR, Rouge-L, and BLEU-1.
+KEANet is supported by additional knowledge that it can generate text descriptions with entity names.
+Therefore, the KEANet outperforms other models in all metrics.
+Fig.
+11 shows the qualitative results on Goal dataset, including generation results of V2C, Clip4Caption, SwinBert and our KEANet.
+Due to the lack of relevant tokenizer tools in V2C, it cannot process text with entity names and can only use special symbols to replace specific entity names.
+Compared to the other three models, KEANet can generate correct entity names and partial actions.
+The above results indicate that, even in challenging tasks such as sports commentary, with the support of additional knowledge, the model can generate text with entity names.
+TABLE VI: Ablation study on MbgVC.
 Ki and Kn denote the player images knowledge and player names knowledge, respectively.
-Attn.
-is the multi-head self-attention module.
+Ea.
+is the entity-aware module.
 Bi-GRU is bi-directional GRU module.
 Numbers in bold indicate the best performance.
-5.3 Ablation Study
-To verify the contribution of the external knowledge and other modules in EAC model, we also perform an ablation study by progressively adding these as input.
-The results of ablation study are shown in Tab.
-4.
+Model Ki Kn Ea. Bi-GRU CIDEr METEOR Rouge-L BLEU-4 RoleF1
+20.1 20.6 40.9 23.0 3.4
+23.5 22.7 43.1 24.4 3.6
+18.0 20.0 38.0 23.2 3.3
+110.6 25.2 49.1 27.3 12.4
+115.7 25.6 53.1 28.1 14.8
+119.3 26.8 53.3 29.5 16.5
+122.2 27.3 54.6 30.5 17.1
+138.5 28.0 54.9 32.4 20.6
+Figure 10: Data sample from the revised Goal dataset.
+Each video is annotated by fileid, team_id, caption, player images and player names.
+Each of the players involved in caption as well as their teammates serve as candidate players.
+The caption in this sample includes the names of players and team.
+Figure 11: Qualitative results on Goal dataset.
+(V2C: Video2Commonsense; C4C: Clip4Caption; SwB: SwinBert; KEA: our proposed model; GT: the ground truth).
+The different specific entity names are marked in red.
+And the fine-grained actions are marked green.
+Since V2C does not have its own tokenizer and vocabulary list, it cannot decode names.
+So names are replaced with the special token $<$UNK$>$.
+V-E Ablation Study
+To verify the contribution of the additional knowledge and other modules in KEANet, we also perform an ablation study by progressively adding these as input.
+The results of ablation study are shown in Table VI.
 Model ① consists of Resnet-18 and T5 decoder, with its input being solely video.
-From the comparison results between of model ① and model ②, it can be observed that adding players’ images to model based on ① does not improve the model’s performance.
+Model ② adds the Bi-GRU module on top of Model ①.
+The Bi-GRU models the temporal contextual relationships between video frames, enabling the model to better understand the dynamic information in videos and predict more accurate action categories.
+Therefore, Model ② has shown improvements in all metrics relative to Model ①, except for the RoleF1 score.
+From the comparison results between of model ① and model ③, it can be observed that adding players’ images to model based on ① does not improve the model’s performance.
 During the training stage, if the model has not been exposed to enough specific names, or has not learned how to infer names from video and image features, it may not be able to generate these names.
-The comparison results between model ① and model ③ highlight a significant improvement by solely adding players’ names to the model based on ①.
+The comparison results between model ① and model ④ highlight a significant improvement by solely adding players’ names to the model based on ①.
 Providing a list of names as input allows the model to effectively integrate this information with the video features, leading to more accurate generation of names.
-This is likely providing the model with external contextual information to help it make more accurate predictions.
-From the comparison results between model ③ and model ④, we can find that player images can emphasize the roles based on entity names and improve the names’ prediction.
-From the comparison results of models ④-⑦, it can be observed that adding multi-head self-attention and Bi-GRU modules can improve the performance of the model.
-The multi-head self-attention module can model the associations among players and focus on key players.
-The Bi-GRU module can model the temporal information of video frame features, capturing the action information.
-From the above results, the external players knowledge brings significant gains in entity awareness.
-The model can generate text descriptions with specific player names and fine-grained actions through external knowledge and temporal modeling.
-5.4 Error Analysis
-The error analysis are conducted by presenting cases of EAC.
+This is likely providing the model with additional context information to help it make more accurate predictions.
+From the comparison results between model ④ and model ⑤, we can find that player images can emphasize the roles based on entity names and improve the names’ prediction.
+From the comparison results of models ⑤-⑧, it can be observed that adding entity-aware and Bi-GRU modules can improve the performance of the model.
+The entity-aware module can model the associations among players and focus on key players.
+The Bi-GRU module can model the temporal contextual information of video frame features, capturing the action information.
+From the above results, the additional players knowledge brings significant gains in entity awareness.
+The model can generate text descriptions with specific player names and fine-grained actions through additional knowledge and temporal modeling.
+Figure 12: Representative error cases of the generated captions, which correspond to the player mismatching, action confusion and lack distance-aware perception.
+GT denotes the ground truth description and PR denotes the generated description.
+V-F Error Analysis
+The error analysis are conducted by presenting cases of KEANet on proposed VC_NBA_2022 dataset.
 As shown in Fig.
-5, we show several errors.
+12, we show several errors.
 (1) Player Mismatching: when the number of players in the sentence is more than one, the model may decode the name of one of the players incorrectly.
@@
 For example, the player appears to be shooting a three-point shot, but is actually inside the three-point line.
-Figure 5: Representative error cases of the generated captions, which correspond to the player mismatching, action confusion and lack distance-aware perception.
-GT denotes the ground truth description and PR denotes the generated description.
-Given the above result, we discuss some potential ways for developing an advanced models for KABVC task.
+Given the above result, we discuss some potential ways for developing an advanced models for KEBVC task.
 First, beyond the current object detection, it is necessary to enhance the model’s ability to understand and handle names.
@@
 This indicates that we need to further improve the spatial awareness of our model.
-6 Conclusion
-In this paper, we introduce the task of knowledge aided basketball video captioning (KABVC) for live text broadcast.
-To achieve this task, we further construct a multimodal basketball game knowledge graph (MbgKG) to provide the external knowledge.
+VI Conclusion
+In this paper, we introduce the task of knowledge guided entity-aware basketball video captioning (KEBVC) for live text broadcast.
+To achieve this task, we further construct a multimodal basketball game knowledge graph (KG_NBA_2022) to provide the additional knowledge.
 Through partial relationships in the knowledge graph, a new multimodal basketball game video captioning dataset is then constructed.
-We also propose a new metric GDS for video captioning, which consider the accuracy of entity name prediction.
-Furthermore, our experiments validate the importance of additional knowledge, including player images and entity names, for basketball video captioning.
-Appendix A Construction of MbgKG
-Figure 6: An example of the play-by-play page.
-Figure 7: Examples of 9 types of basketball events. Each type has the corresponding keyword marked in red.
-Figure 8: Examples of player information and team information.
-Figure 9: An example of the multi-tuples parsed from event descriptions, players’ information, and teams’ information.
-In this multi-tuples, the home team information is marked in red, the away team information is marked in blue, and events are marked in green.
-The example of play-by-play page is shown in Fig.
-6.
-This page not only contains the event descriptions, but also contains links to player information and team information.
-We can collect relevant players’ information and teams’ information through these links.
-All the collected event descriptions can be divided into 9 categories, as shown in Fig.
-7.
-Each event has its own unique keywords which are marked in red.
-In addition to keywords, there are also some special words in different events.
-For example, in shooting events, there are words such as “assist by” and “block by”.
-By obtaining the indices of these keywords and special words, we can parse the descriptions into multi-tuples.
-Fig.
-8 shows the players’ information and teams’ information.
-We only collect information that is available for most players, including “name”, “position”, “shoots”, “height”, “weight”, “birth”, “NBA debut”, and the link to their profile pictures.
-For teams’ information, we only collect “name”, “season”, “coach”, “executive”, “arena”, and the link to teams’ logos.
-By using OCR technology to recognize the time on the scoreboard in each video frame, we can compare it to the time on the play-by-play page.
-If the recognized time matches the time on the scoreboard, the video timestamp of the current event is recorded in the multi-tuples.
-Fig.
-9 shows the example of the multi-tuples parsed from event descriptions, players’ information, and teams’ information.
-Through these efforts, a Multimodal Basketball Game Knowledge Graph (MbgKG) containing 11489 events and 42870 relationships is constructed.
-MbgKG not only contains the static information, but also includes the temporal dynamic information.
-We believe that this knowledge graph will continue to be extended and be explored more valuable knowledge for sports video understanding.
-Appendix B Construction of MbgVC
-In basketball, shot and rebound events are the most prevalent.
-Meanwhile, we count the number of different types of events in MbgKG.
-Tab.
-5 shows that there are 4400 shot events and 2649 rebound events, which are the most common types of events.
-Based on the aforementioned common knowledge and statistical data, videos and event descriptions about shot and rebound are extracted from MbgKG to construct the dataset.
-Shots can be divided into three types: two-point jump shot (2p jump shot), two-point layup (2pt-layup), and three-point jump shot (3p jump shot).
-Rebounds can be divided into two types: defensive rebound and offensive rebound.
-In addition, scoring and missing a shot are represented by “make” and “miss”, respectively.
-When a shot is not scored, it is often followed by a rebound event.
-Therefore, we combine the missed shot event and rebound event into one event.
-In total, there are nine types of shooting events, which are “2p-succ.”, “2p-fail.-off.”, “2p-fail.-def.”, “2p-layup-succ.”, “2p-layup-fail.-off.”, “2p-layup-fail.-def.”, “3p-succ.”, “3p-fail.-off.” and “3p-fail.-def.” Subsequently, based on these nine types of shooting events and the relationships within MbgKG, we extract videos, text descriptions, and player information from MbgKG.
-It’s worth noting that there are some issues with the text descriptions in MbgKG: 1) The names of the players are abbreviated, which does not meet the requirements of actual applications; 2) The texts contain descriptions of shooting distances, such as ”from xx ft” and ”at rim”.
-This makes it difficult for the captioning model to generate accurate distances; 3) The text descriptions do not conform to the structural and compositional rules of natural language.
-Figure 10: Qualitative results on MbgVC dataset.
-(V2C: Video2Commonsense; C4C: Clip4Caption; SwB: SwinBert; EAC: our proposed model; GT: the ground truth).
-The different specific entity names are marked in red and blue, respectively.
-And the fine-grained actions are marked green.
-Since V2C does not have its own tokenizer and vocabulary list, it cannot decode names.
-So names are replaced with the special token $<$UNK$>$.
-In response to the above issues, we design some modification rules to address them.
-For names, we replace abbreviated names in the text descriptions with the full names from the player attribute information in MbgKG (“L.
-James” $\Longrightarrow$ “LeBron James”).
-The distance descriptions (“from XX ft” and “at rim”) in the texts are directly removed.
-Sometimes, when a player scores a shot, it is assisted by his teammate.
-We modify “(assist by SOMEONE)” to “with an assist from SOMEONE”.
-When a player doesn’t score a shot, sometimes it’s because the ball is blocked by the opposing player.
-We modify “(block by SOMEONE)” to “blocked by SOMEONE”.
-Events Foul Rebound Violation Timeout Freethrow Entergame Turnover Jumpball Shot
-num. 986 2649 68 271 1103 1189 774 49 4400
-Table 5: Statistics of the events in MbgKG. “num.” is the abbreviation of “number”. Numbers in bold indicate the highest two values.
-Appendix C Evaluation Metric
-Our task is oriented towards basketball live text broadcast, where concise and succinct text descriptions can allow users to quickly grasp and understand the content of the game in a short time.
-The names of the players and their corresponding fine-grained actions are the main components of the text descriptions.
-Existing evaluation metrics based on n-gram such as CIDEr, METEOR, Rouge-L, and BLEU, primarily penalize the semantically correct but textually inconsistent text descriptions.
-Each evaluation metric has its unique calculation method and principle, and no single metric can be considered as authoritative and comprehensive.
-Therefore, the most direct approach is to evaluate the quality of the text description by combining multiple metrics.
-Since existing evaluation metrics cannot assess the accuracy of player name prediction, we introduce $e_{acc}$ to address this issue.
-When evaluating concise texts, the effectiveness of Rouge-L diminishes as it relies on the length of the longest common subsequence between the two texts.
-Therefore, compared to other evaluation metrics, the importance of Rouge-L would be somewhat lessened.
-We believe that the assessment of player name prediction is crucial, as it is a fundamental requirement in real-world applications.
-Based on the aforementioned analysis, we assign the highest coefficient to $e_{acc}$ and the lowest coefficient to Rouge-L.
-The coefficients of CIDEr and METEOR fall between those of $e_{acc}$ and Rouge-L.
-From experience, the coefficient of eacc is set to 4, the coefficients of CIDEr and METEOR are set to 2, and the coefficient of Rouge-L is set to 1.
-We sought opinions from several experts and scholars in the field of video captioning, all of whom expressed that the coefficient allocation has a certain degree of rationality.
-Appendix D Qualitative Results
-Fig.
-10 shows the qualitative results on MbgVC dataset, including generation results of V2C, Clip4Caption, SwinBert and our EAC model.
-V2C, Clip4Caption and SwinBert can correctly generate some actions but fail to generate correct entity names because these names never appear during training.
-However, with the help of the external knowledge, our EAC could well relate entities in video clips with the fine-grained actions.
-These cases demonstrate that our newly proposed EAC model more consistently with the requirement of practical application.
+Furthermore, our experiments validate the importance of additional knowledge, including player images and entity names, for entity-aware video captioning.